OpenStudy (stormoshawty):

Simplify the equation. 5x^-4

OpenStudy (anonymous):

x^-n=1/x^n

OpenStudy (stormoshawty):

Where do you get n?

OpenStudy (anonymous):

this is just a proof expression. it explains how to solve your problem

OpenStudy (stormoshawty):

Ohhhh, I was so confused.

OpenStudy (anonymous):

x^-2=1/x^2 does that make more sense?

OpenStudy (stormoshawty):

Yes, but is that the expression simplified?

OpenStudy (anonymous):

it is not, there was a warning the other day about posting answers. so i will help you understand why the answer is what it is.

OpenStudy (stormoshawty):

Alrighty.

OpenStudy (anonymous):

alright so where are you, in solving this problem knowing the reasoning i gave you can you figure it out?

OpenStudy (stormoshawty):

is it 5/x^-4

OpenStudy (anonymous):

your close but remember what happens to the sign

OpenStudy (anonymous):

the confusion is in the wording there is no such mathematical operation as "simplify" it is the lazy math teacher's way of saying "give me the answer i want" the instructions should be "rewrite using positive exponents"

OpenStudy (anonymous):

to make the exponent positive, take the reciprocal i.e. flip it leave the coefficient on top change the exponent from positive to negative example: 4x^-2 =4/x^2

OpenStudy (anonymous):

then ask you math teacher why on earth he or she thinks 5/x^4 is any "simpler" than 5x^-4 there is nothing "simpler" about it, it is just another way to write it with positive rather than negative exponents